OpenStudy (anonymous):

how do i solve for x? 25^(2x) = 125^(x+2)

myininaya (myininaya):

\[25=5^{some power}\] what is that power?

myininaya (myininaya):

\[125=5^?\] also want to write 125, as 5^?

myininaya (myininaya):

So if we can write both sides having the same base, all we need to do is set the exponents equal

myininaya (myininaya):

\[\text{ what is } 5^2?\] \[\text{ what is } 5^3 ?\]

OpenStudy (anonymous):

5^2 = 25 5^3 = 125

myininaya (myininaya):

yeah so we can do the following : \[(5^2)^{2x}=(5^3)^{x+2}\]

myininaya (myininaya):

\[5^{2 \cdot 2x}=5^{3(x+2)}\]

myininaya (myininaya):

so the bases are the same set the exponents equal to each other and solve for x

OpenStudy (anonymous):

k, thanks

OpenStudy (anonymous):

does x= 6

myininaya (myininaya):

\[2 \cdot 2x=3(x+2)\] \[4x=3x+6\] \[4x-3x=6\] \[1x=6 \] \[x=6\]

myininaya (myininaya):

yep